Home CMS Production Clean room MedAustron HEPHY testbeams old
testbeam-DESY-Oct2017 testbeam-DESY-April2017 testbeam-SPS2015 testbeam-SPS2014 testbeam-DESY14 testbeam-SPS12 testbeam-SPS11 testbeam-SPS10 testbeam-SPS09 testbeam-SPS08 LP-TPC
  HEPHY logbook for testbeam at SPS June 2008  Not logged in ELOG logo
Message ID: 18     Entry time: Sat May 31 17:36:14 2008
Author: Markus Friedl 
Project: SiLC 
Subject: APVDAQ + telescope (0-sup) 
Run Number: run2718 
Events: 100000 
StartTime: 31.05.2008 16:21:57 
EndTime: 31.05.2008 21:08:32 
SiLC beam test - CERN SPS June 2008
8 modules, HV=100V, 40MHz, 6 samples, hor stage :=+2 mm
trigger: 9 mm opening, TLUControl.exe -d 10 -a 12 (now for real)
telescope: zero supressed, new firmware

- long break in the middle due to power outage after thunderstorm

horizontal position ("target") = +2.0mm - after we decided to scan the whole area in 3 steps
this is RIGHT position

all 3 positions are:
RIGHT: +2.00 mm    <--- this run
MID:   -2.75 mm
LEFT:  -7.50 mm
ELOG V3.1.4-966e3dd